A sample that contains 2.54 mol of a gas at 298 K has a pressure of 2.16 atm. What is its volume?
1
Expert's answer
2020-04-13T02:28:08-0400

According to the equation of Mendeleev-Clapeyron for gases ("p•V = \nu•R•T" ), we can determine the volume using this formula (considering that 2.16 atm = 202650 Pa, Pa = N/m2, J = N•m):

"V = (\nu•R•T) / p" = (2.54(mol)•8.31((N•m)/(mol•K))•298(K)) / 202650(N/m2) = 0.03 m3 = 30 ml

Answer: V=30 ml